Circular No. 8903

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 8P/TUTTLE D. Schleicher, Lowell Observatory; and L. Woodney, California State

University, San Bernardino, write that they obtained CN narrowband imaging of comet 8P on Dec. 14 using the Hall 1.1-m telescope at Lowell Observatory. Following the removal of a mean radial profile, a series of three successive spiral arcs are seen, approximately centered in the sunward direction and extending more than 180^{o}. Both the spacing and the outward motion of the arcs during a 4-hr interval imply a rotation period of 4.9–5.0 hr, assuming that a single jet is responsible for all of the arcs. The projected CN gas velocity towards the position angle of the sun is ∼ 1.1 km/s.

Schleicher adds that he obtained fourteen sets of narrowband photometry of comet 8P on Dec. 3, 4, and 5 (when r = 1.31–1.29 AU) using the Hall 1.1-m telescope at Lowell Observatory, resulting in the following averaged production rates: log Q(OH) = 27.80; equivalent log Q(water; vectorial) = 27.88; log Q(NH) = 25.64; log Q(CN) = 25.22; log Q(C_{2}) = 25.39; log Q(C_{3}) = 24.82; log Afρ(526.0 nm) = 1.5 (cf. IAUC 7342). Dust Afρ exhibits a strong increasing trend with decreasing aperture size that is due, at least in part, to a non-negligible nucleus signal.

Total visual magnitude estimates by J. J. Gonzalez, Leon, Spain: Nov. 3.85 UT, 12.1 (0.20-m reflector); 9.92, 11.3 (25×100 binoculars); 13.85, 10.5; 27.79, 9.3; Dec. 5.01, 8.6 (10×50 binoculars); 11.79, 7.6.

COMET P/2006 W4 = P/1993 D1 (HILL) Corrigendum. Regarding the four astrometric positions published on

IAUC 8902, the Feb. 26 data are from 1993, and the Apr. 5 observations are from 1994.

VARIABLE STAR IN SERPENS E. O. Waagen, AAVSO, and H. Yamaoka, Kyushu University, note

that the variable star reported on IAUC 8898 appears to be a very bright IRAS source (IRAS 18066–0722). Yamaoka adds that it also appears to be catalogued by the 2MASS project, and Waagen adds that the variable apparently is listed by the ASAS3 project as a variable varying from V = 13.7 to [15 with a light curve that is suggestive of a Mira-type variable with maximum around Dec. 5. Yamaoka also reports that a low-resolution spectrum was taken by K. Kinugasa and O. Hashimoto, Gunma Astronomical Observatory (GAO), on Nov. 24.35 UT with the GAO 1.5-m telescope, showing that this object is a very red star — likely a Mira-type variable.

2007 December 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green